Hello,
I just had this question awnsered, but i am confused.
The question was 1/p-1-q=1/f solve for f
I was told;
1. It will simplify things a bit if we first subtract the fractions on the left side.
and got this:
q-p/pq=1/f

I am confused how you got that! please help!Please show step and tell why.
Thank you


Answer by duggie(26)

Understand now?


Simplifying
1p + -1 + -1q = 1f
Reorder the terms:
-1 + 1p + -1q = 1f
Solving
-1 + 1p + -1q = 1f
Solving for variable 'p'.
Move all terms containing p to the left, all other terms to the right.
Add '1' to each side of the equation.
-1 + 1p + 1 + -1q = 1 + 1f
Reorder the terms:
-1 + 1 + 1p + -1q = 1 + 1f
Combine like terms: -1 + 1 = 0
0 + 1p + -1q = 1 + 1f
1p + -1q = 1 + 1f
Add 'q' to each side of the equation.
1p + -1q + q = 1 + 1f + q
Combine like terms: -1q + q = 0
1p + 0 = 1 + 1f + q
1p = 1 + 1f + q
Divide each side by '1'.
p = 1 + 1f + q
Simplifying
p = 1 + 1f + q